# Supersymmetric Black Hole Hair and $AdS_{3}\times S^{3}$

### v1

@@ Abstract
The 4D-5D connection allows us to view the same near horizon geometry as part of a 4D black hole or a 5D black hole. A much studied example of this phenomenon is the BMPV black hole uplifted to 6D with flat base space versus Taub-NUT base space. These black holes have identical near horizon $AdS_{3}\times S^{3}$ geometry. In this paper, we study modes in $AdS_{3}\times S^{3}$ and identify those that correspond to supersymmetric hair modes in the full black hole spacetimes. We show that these modes satisfy non-normalisable boundary conditions in $AdS_{3}$. The non-normalisable boundary conditions are different for different hair modes. We also discuss how the supersymmetric hair modes on BMPV black holes fit into the classification of supersymmetric solutions of 6D supergravity.
@@ 1 Introduction
In section 4, we study the deformation due to turning on an anti-self-dual three-form field strength in $AdS_{3}\times S^{3}$. Using the decomposition of the anti-self-duality equation in spherical harmonics, following [13], we derive equations for such perturbations in $AdS_{3}$. Solving these equations, we identify the anti-self-dual field strength that correspond to a hair mode of the original black hole. Then, solving the Einstein equations in $AdS_{3}\times S^{3}$ with this anti-self-dual field as an additional source, we obtain the deformed metric. We do a regularity analysis for this mode, and find that both the metric and the anti-self dual field strength are regular at the $AdS_{3}$ horizon. As for the case of the GV modes, the anti-self-dual field strength deformations are also not asymptotically $AdS_{3}\times S^{3}$ but correspond to non-normalizable deformations even though in the full black hole geometry they are normalizable. Moreover, we find that the non-normalizable metric deformation in this case is closely related to null-warped $AdS_{3}$.
@@ 2 Review: black holes, near horizon geometry, and bosonic hair modes
with $r_{0}$ related to the charges carried by the black hole. The coordinate $r$ is the Gibbons-Hawking radial coordinate on 4D flat space introduced as follows. A standard set of coordinates for four-dimensional Euclidean flat space are
@@
related to cartesian coordinates as,
@@
In these coordinates the flat space metric takes the form
@@
Next we consider $J_{L}=0$ BMPV black hole in Taub-NUT space [14, 15], following the notation of refs. [5, 7]. The metric of the four dimensional Taub-NUT space in Gibbons-Hawking coordinates is given by
@@
Now we consider the construction of hair modes [5, 7]. The solution-generating technique of Garfinkle and Vachaspati [18, 19] was used to add hair modes to the above systems. Given a space-time metric $G_{MN}$ with null Killing vector $k_{M}$ that is hypersurface orthogonal, i.e.,
@@
for some function $A$, a new exact solutions to the supergravity equations are constructed by the following transform,
@@
with $e^{A}=\psi$. Applying the Garfinkle-Vachaspati transform we get,
@@
with four arbitrary functions $f_{i}(v)$. The deformed metric (2.26) does not look asymptotically flat, but via a standard change of coordinates [20] it can be brought to a manifestly asymptotically flat form.
@@
Next we recall the hair mode deformations of the BMPV black hole in Taub-NUT space. A class of these deformations is generated by Garfinkle-Vachaspati transform. The metric for these deformations take the form
@@
where $y^{i}$ are cartesian coordinates on $\mathbb{R}^{3}$ and $g_{i}(v)$ are three arbitrary functions.
The BMPV black hole in Taub-NUT space admits another class of deformations [5] sourced by anti-self-dual three-forms. In this paper we consider only one such three-form to be present. These new deformations arise (essentially) because the Taub-NUT space admits a self-dual harmonic form,
@@ 3 Garfinkle-Vachaspati hair modes in the near-horizon region
where $P^{(a,b)}_{n}$ are the Jacobi polynomials and $x=\cos 2\tilde{\theta},S=m_{+}+m_{-},D=m_{+}-m_{-}$. We will see below that the black hole hair modes discussed in the previous section correspond to different $L$ modes from the $AdS_{3}\times S^{3}$ perspective. In $(\bar{u},\bar{v},\bar{z})$ coordinates, the deformed metric (3.1) takes the form
@@ 3.1 $L=0$ modes
With $L=0$ equation (3.5) simply reduces to the wave equation in $AdS_{3}$. Since there are no propagating degrees of freedom in three-dimensional gravity, we expect for (3.4) either a trivial solution or a solution which is related to $AdS_{3}\times S^{3}$ by a coordinate transformation. Indeed, this expectation is realised. Solving equation (3.2), we get (the index $I$ takes the value $(0,0,0)$ and is dropped for simplicity),
@@ 3.2 $L=1$ modes
It is clear that the normalisable modes $c^{I}(v)z^{3}$ are not regular at the horizon. The non-normalisable modes $\frac{d^{I}(v)}{z}$ can be regular at the horizon. The index $I$ takes values $(1,m_{+},m_{-})$ with $m_{+}=\pm\frac{1}{2},m_{-}=\pm\frac{1}{2}$. Discarding the normalisable modes, we get,
@@
These additional terms all vanish in the $V\to 0$ limit. The resulting metric is thus smooth near $V=0$, however, the $V$ derivatives of the metric are not. This leads to divergences in the Riemann tensor. For example, for $R_{VWVW}$ diverges as $1/V$ as $V\to 0$. The conclusion is in line with the fact that these modes are also singular on the BMPV black hole horizon in the full black hole geometry [5].
@@ 3.3 $L=2$ modes
We see that the function $H$ in (3.25) corresponds to $g_{i}(v)y^{i}$ for $i=1,2,3$ and hence to the Garfinkle-Vachaspati hair modes for the black hole in Taub-NUT space (2.29).
@@
$ds^{2}=\ell^{2}\left(W^{2}dUdV+\frac{dW^{2}}{W^{2}}+W^{4}g_{i}(\bar{v}(V))n^{i}dV^{2}\right)+\ell^{2}d\Omega^{2}_{S^{3}}$ (3.27)
@@
We conclude that the three functions $g_{i}(v)$ generate smooth deformations of the near-horizon geometry. [^3] We can now qualify the statement in [22] that $L=2$ modes are not regular in bulk. We can pick $L=2$ solutions which are, in fact, regular in the bulk but are non-normalisable. This conclusion is in line with the analysis of [5] where it was shown that the three Garfinkle-Vachaspati hair modes for the black hole in Taub-NUT space are also regular at the horizon.
@@ 4 Form field hair modes in the near horizon region
We now consider a deformation of $AdS_{3}\times S^{3}$ due to the anti-self-dual form field $H_{MNP}$. The relevant perturbation equations were given in [13]. For our analysis the most relevant equation is the anti-self-duality condition for the perturbing form field with mixed components. This equation takes the form [13, eq. (4.18)]
@@
with $\eta=-2$. For later use we keep $\eta$ general. To solve equation (4.7) we make the ansatz that only $b_{v}$ component of $b_{\rho}$ is non-vanishing. [^4] Applying $\nabla^{\mu}$ on (4.7) we see that $\nabla^{\mu}b_{\mu}=0$, where $\nabla^{\mu}$ is the covariant derivative compatible with the three-dimensional $AdS_{3}$ part of the metric. We can use this condition to set $b_{z}=0$. Next, requiring $db$ to be compatible with the $k^{M}$ Killing symmetry, we have £$\pounds_{k}db=d(i_{k}db)=0$. Choosing $i_{k}db=0$ gives us $i_{k}(\star_{3}b)=0$ from (4.7), which implies $\star_{3}(b\wedge k)=0$. This condition can be used to set $b_{u}=0$. Furthermore, we assume that the perturbation is compatible with $\partial_{u}$ Killing symmetry, i.e., $b_{v}$ only depends on the $v$ and $z$ coordinates. Substituting this ansatz in equation (4.7), we get
@@
This simple perturbation describes the anti-self-dual form-field perturbations to the black hole in Taub-NUT space from the near horizon region perspective.
To see this, let us recall that the three-form hair perturbation for the black hole in Taub-NUT space takes the form (2.31). In the near horizon limit $r\to 0$ the $H$-field takes the form,
$H=-\frac{h(v)r}{R_{4}^{2}}\sin\theta\,dv\wedge d\theta\wedge d\phi+\frac{h(v)}{R_{4}^{2}}dv\wedge dr\wedge(dx^{4}+\cos\theta d\phi).$ (4.11)
In coordinates $(u,v,z,\tilde{\theta},\tilde{\phi},\tilde{\psi})$, this becomes
@@
Now that we have a solution to the anti-self dual form-field, we can consider its back reaction on the metric and solve the corresponding Einstein equation with the stress tensor sourced by the form-field deformation. In our conventions, Einstein equation with anti-self-dual form field as source an additional source is [5, 7]
@@
With the above $H_{MNP}$ (4.12)–(4.13) a simple calculation shows that it only sources the $vv$ component of the Einstein equation, and moreover the sphere dependence drops out. We find [^5] It is simplest to write this result in mixed component. We will see shortly that the left hand side of the Einsteins equations take a particularly simple form with mixed indices of this type.
@@
This is precisely the $r\to 0$ limit of the function $S$ that appears in equation (2.33). We also note that the three-dimensional part of the metric (4.16) is of the following special form,
@@
$ds^{2}=\ell^{2}\left(W^{2}dUdV+\frac{dW^{2}}{W^{2}}-W^{4}\bar{h}(\bar{v}(V))^{2}dV^{2}\right)+\ell^{2}d\Omega^{2}_{S^{3}}$ (4.20)
@@ 5 Fermionic hair modes in the near horizon region
In this section we make some observations on the non-normalisability of the fermionic hair modes of refs. [5, 4, 7] in $AdS_{3}\times S^{3}$. We find it easiest to present this analysis in six-dimensions. The relevant linearised equations for the gravitino $\psi^{\alpha}_{M}$ are
@@ 6 Discussion
    Using the AdS/CFT dictionary, we can identify the dimensions of the CFT operators that correspond to the bulk deformations discussed above. Turning on non-normalisable modes correspond to deformations of the CFT with operators of conformal dimension $\Delta>2$. For operators dual to scalars of mass $m$ in $AdS_{3}$, $\Delta=1+\sqrt{1+m^{2}\ell^{2}}$. The Garfinkle-Vachaspati deformations in $AdS_{3}\times S^{3}$ with $L=1$ and $L=2$ can be identified with scalars of masses $m^{2}\ell^{2}=L(L+2)$. Hence, the smooth hair mode with $L=2$ corresponds to a deformation of the CFT by an operator of mass dimensions $\Delta=4$. The same conclusion can be reached by considering the Garfinkle-Vachaspati deformation as a metric deformation. Since the deformation term grows as $z^{-4}$ near the boundary $z=0$, it corresponds to a deformation of the CFT by an operator of mass dimensions $\Delta=4$, see, e.g., comments in [23].
@@
    We can similarly identify the dimensions of the operators that correspond to the form-field deformation. Taking one more exterior derivative of equation (4.7), we can convert it to the equation for a massive vector field
@@
    In section 4, we made the observation that the three-dimensional part of the metric (4.16) can be written in a special form (4.19). For the case when the function $h(v)$ is a constant, we recognise metric (4.19) as the null warped $AdS_{3}$ metric (and also as Schrödinger metric with $z=2$). As is well known, these metrics can be obtained as solutions to three-dimensional Einstein gravity with negative cosmological constant coupled to a massive vector field [24]. As discussed in the previous remark, the metric deformation discussed in section 4 can also be thought of as a deformation sourced by a massive vector field (6.1). Thus, the fact that the metric (4.16) can be written in the null warped $AdS_{3}$ form is perhaps not unexpected. Null warped $AdS_{3}$ metrics are known to be duals of deformations of the boundary CFT by operators of dimension $4$ [25]. This is in line with the our observations made above. It will be interesting to identify such operators in the D1-D5 CFT and explore the connection to the null warped $AdS_{3}$ further.
In refs. [26, 27, 28] various deformations of $AdS_{3}\times S^{3}$ were considered in the context of the fuzzball program. Even though the starting point is the same, there are many differences between our results and their analyses. Firstly, since there are no horizons in the fuzzball cases, these references did not need to consider the regularity at the horizon. The analyses in these references is instead in global AdS. Secondly, the metric perturbations considered in these references do not have direct analogs in our analysis, but roughly speaking a class of modes considered in [26] corresponds to our $L=0$ Garfinkle-Vachaspati modes. See also point 3 on page 26 of ref. [26] where a related observation is made. In [29], it was suggested that the fuzzballs should also admit hair modes like the ones that we have considered in this paper. Specifically, by putting fuzzballs in Taub-NUT space, it was conjectured that there should be fuzzball hair, like the ones obtained by putting the BMPV black hole in Taub-NUT space. We hope to report our progress on constructing such hair in the future.

### v2

@@ Abstract
The 4D-5D connection allows us to view the same near horizon geometry as part of a 4D black hole or a 5D black hole. A much studied example of this phenomenon is the BMPV black hole uplifted to 6D with flat base space versus Taub-NUT base space. These black holes have identical near horizon $AdS_{3}\times S^{3}$ geometry. In this paper, we study modes in $AdS_{3}\times S^{3}$ and identify those that correspond to supersymmetric hair modes in the full black hole spacetimes. We show that these modes satisfy non-normalisable boundary conditions in $AdS_{3}$. The non-normalisable boundary conditions are different for different hair modes and for different asymptotic completion. We also discuss how the supersymmetric hair modes on BMPV black holes fit into the classification of supersymmetric solutions of 6D supergravity.
@@ 1 Introduction
At the linearised level, it was already understood in ref. [4] that hair modes are located outside the near-horizon region. The analysis of this paper makes those observations more quantitative. It is useful to appreciate another context where a closely related issue is well explored, namely, quasi-normal modes of microstate geometries [13, 14, 15, 16]. In this literature, it is well appreciated that in the limit when a microstate geometry has a large $AdS_{3}\times S^{3}$ region, quasi-normal modes correspond to excitations that leak out from the near-core $AdS_{3}\times S^{3}$ region to flat space region. The leaking excitations satisfy non-normalisable boundary condition at the edge of $AdS_{3}$ (otherwise they won’t leak out). Our results are exactly the same, but in the context of the supersymmetric hair modes. The hair mode “leak out” from the near-throat region. The hair modes are supported outside the throat region and they decay to zero in the asymptotic region.
In section 4, we study the deformation due to turning on an anti-self-dual three-form field strength in $AdS_{3}\times S^{3}$. Using the decomposition of the anti-self-duality equation in spherical harmonics, following [17], we derive equations for such perturbations in $AdS_{3}$. Solving these equations, we identify the anti-self-dual field strength that correspond to a hair mode of the original black hole in Taub-NUT space. Then, solving the Einstein equations in $AdS_{3}\times S^{3}$ with this anti-self-dual field as an additional source, we obtain the deformed metric. We do a regularity analysis for this mode, and find that both the metric and the anti-self dual field strength are regular at the $AdS_{3}$ horizon. As for the case of the GV modes, the anti-self-dual field strength deformations are also not asymptotically $AdS_{3}\times S^{3}$ but correspond to non-normalizable deformations even though in the full black hole geometry they are normalizable. Moreover, we find that the non-normalizable metric deformation in this case is closely related to null-warped $AdS_{3}$.
@@ 2 Review: black holes, near horizon geometry, and bosonic hair modes
with $r_{0}$ related to the charges carried by the black hole. The coordinate $r$ is the Gibbons-Hawking radial coordinate on 4D flat space introduced as follows. The spherical polar coordinates for four-dimensional Euclidean flat space $(\tilde{r},\tilde{\theta},\tilde{\phi},\tilde{\psi})$, in which the metric takes the form,
@@
are related to cartesian coordinates as,
@@
In these coordinates flat space metric takes the form
@@
Next we consider $J_{L}=0$ BMPV black hole in Taub-NUT space [18, 19], following the notation of refs. [5, 7]. The metric of the four dimensional Taub-NUT space in Gibbons-Hawking coordinates is given as a U(1) fibre over an $\mathbb{R}^{3}$ base space,
@@
Now we consider the hair modes [5, 7]. The solution-generating technique of Garfinkle and Vachaspati [22, 23] was used to add hair modes to the above systems. Given a space-time metric $G_{MN}$ with null Killing vector $k_{M}$ that is hypersurface orthogonal, i.e.,
@@
for some function $A$, new exact solutions to the supergravity equations are constructed by the following transform,
@@
with
$e^{A}=\psi.$ (2.26)
Applying the Garfinkle-Vachaspati transform we get,
@@
with four arbitrary functions $f_{i}(v)$. The deformed metric (2.27) does not look asymptotically flat, but via a standard change of coordinates [24] it can be brought to a manifestly asymptotically flat form. We can take the near-horizon limit of the deformed metric by taking the $r\rightarrow 0$ limit. It gives
$ds^{2}=\frac{\ell^{2}}{z^{2}}\left(dudv+dz^{2}+\left(\frac{z^{2}}{\ell^{2}}+T(z)\right)dv^{2}\right)+\ell^{2}d\Omega^{2}_{S^{3}}.$ (2.29)
where
$T(z)=f_{i}(v)w^{i}(z).$ (2.30)
@@
Next we recall the hair mode deformations of the BMPV black hole in Taub-NUT space. A class of these deformations is generated by the Garfinkle-Vachaspati transform. The deformed metric takes the form
@@
where $y^{i}$ are cartesian coordinates on $\mathbb{R}^{3}$ and $g_{i}(v)$ are three arbitrary functions. As before, we can take the near-horizon limit of the deformed metric by taking the $r\rightarrow 0$ limit.
The BMPV black hole in Taub-NUT space admits another class of deformations [5] sourced by anti-self-dual three-forms. To keep things as simple as possible, in this paper, we consider only one such three-form to be present. These new deformations arise (essentially) because the Taub-NUT space admits a self-dual harmonic form,
@@
Once again, we can take the near-horizon limit of the deformed metric by taking the $r\rightarrow 0$ limit.
@@ 3 Garfinkle-Vachaspati hair modes in the near-horizon region
where $P^{(a,b)}_{n}$ are the Jacobi polynomials and $x=\cos 2\tilde{\theta},S=m_{+}+m_{-},D=m_{+}-m_{-}$. We will see below that the black hole hair modes discussed in the previous section correspond to different $L$ modes from the $AdS_{3}\times S^{3}$ perspective. In Poincaré coordinates $(\bar{u},\bar{v},\bar{z})$ introduced in (2.20), the deformed metric (3.1) takes the form
@@ 3.1 $L=0$ modes
With $L=0$, equation (3.5) simply reduces to the wave equation in $AdS_{3}$. Since there are no propagating degrees of freedom in three-dimensional $AdS_{3}$ gravity, we expect for (3.4) either a trivial solution or a solution which is related to $AdS_{3}\times S^{3}$ by a coordinate transformation. Indeed, this expectation is realised. Solving equation (3.2), we get (the index $I$ takes the value $(0,0,0)$ and is dropped for simplicity),
@@ 3.2 $L=1$ modes
The normalisable modes $c^{I}(v)z^{3}$ are not regular at the horizon $z\to\infty$. The non-normalisable modes $\frac{d^{I}(v)}{z}$ can be regular at the horizon. The index $I$ takes values $(1,m_{+},m_{-})$ with $m_{+}=\pm\frac{1}{2},m_{-}=\pm\frac{1}{2}$. Discarding the normalisable modes, we get,
@@
These additional terms all vanish in the $V\to 0$ limit. The resulting metric is thus smooth near $V=0$, however, the $V$ derivatives of the metric are not. This leads to divergences in the Riemann tensor. For example, $R_{VWVW}$ diverges as $1/V$ as $V\to 0$. The conclusion is in line with the fact that these modes are also singular on the BMPV black hole horizon in the full black hole geometry [5].
@@ 3.3 $L=2$ modes
We see that the function $H$ in (3.25) corresponds to $g_{i}(v)y^{i}$ for $i=1,2,3$ and hence to the Garfinkle-Vachaspati hair modes for the black hole in Taub-NUT space (2.32). Indeed, by the taking the $r\to 0$ limit of the GV deformed black hole metric in Taub-NUT we obtain $AdS_{3}\times S^{3}$ metric (3.1) with $H$ given as in (3.25).
@@
$ds^{2}=\ell^{2}\left(W^{2}dUdV+\frac{dW^{2}}{W^{2}}+W^{4}g_{i}(\bar{v}(V))n^{i}dV^{2}\right)+\ell^{2}d\Omega^{2}_{S^{3}}.$ (3.27)
@@
We conclude that the three functions $g_{i}(v)$ generate smooth deformations of the near-horizon geometry. [^3] We can now qualify the statement in [26] that $L=2$ modes are not regular in the bulk. We can pick $L=2$ solutions which are, in fact, regular in the bulk but are non-normalisable. This conclusion is in line with the analysis of [5] where it was shown that the three Garfinkle-Vachaspati hair modes for the black hole in Taub-NUT space are also regular at the horizon.
@@ 4 Form field hair modes in the near horizon region
We now consider a deformation of $AdS_{3}\times S^{3}$ due to the anti-self-dual form field $H_{MNP}$. The relevant perturbation equations were given in [17]. For our analysis the most relevant equation is the anti-self-duality condition for the perturbing form field with mixed components. This equation takes the form [17, eq. (4.18)]
@@
with $\eta=-2$. For later use we keep $\eta$ general. To solve equation (4.7) we make the ansatz that only the $b_{v}$ component of the one-form $b_{\mu}$ is non-vanishing. [^4] Applying $\nabla^{\mu}$ on (4.7) we see that $\nabla^{\mu}b_{\mu}=0$, where $\nabla^{\mu}$ is the covariant derivative compatible with the three-dimensional $AdS_{3}$ part of the metric. We can use this condition to set $b_{z}=0$. Next, requiring $db$ to be compatible with the $k^{M}$ Killing symmetry, we have £$\pounds_{k}db=d(i_{k}db)=0$. Choosing $i_{k}db=0$ gives us $i_{k}(\star_{3}b)=0$ from (4.7), which implies $\star_{3}(b\wedge k)=0$. This condition can be used to set $b_{u}=0$. Furthermore, we assume that the perturbation is compatible with $\partial_{u}$ Killing symmetry, i.e., $b_{v}$ only depends on the $v$ and $z$ coordinates. Substituting this ansatz in equation (4.7), we get
@@
As will shortly become clear, this perturbation describes the anti-self-dual form-field perturbations to the black hole in Taub-NUT space from the near horizon region perspective.
To see this, let us recall that the three-form hair perturbation for the black hole in Taub-NUT space takes the form (2.34). In the near horizon limit $r\to 0$ the $H$-field becomes,
$H=-\frac{h(v)r}{R_{4}^{2}}\sin\theta\,dv\wedge d\theta\wedge d\phi+\frac{h(v)}{R_{4}^{2}}dv\wedge dr\wedge(dx^{4}+\cos\theta d\phi),$ (4.11)
which in coordinates $(u,v,z,\tilde{\theta},\tilde{\phi},\tilde{\psi})$, takes the form
@@
Now that we have a solution to the anti-self dual form-field, we can consider its back reaction on the metric and solve the corresponding Einstein equation with the stress tensor sourced by the form-field deformation. In our conventions, Einstein equation with anti-self-dual form field as an additional source take the form [5, 7]
@@
With the $H_{MNP}$ given in (4.12)–(4.13) a simple calculation shows that it only sources the $vv$ component of the Einstein equation, and moreover the sphere dependence drops out. We find [^5] It is simplest to write this result in mixed component. We will see shortly that the left hand side of the Einsteins equations take a particularly simple form with mixed indices of this type.
@@
This is precisely the $r\to 0$ limit of the function $S$ that appears in equation (2.36). We conclude that the B-field (4.10) together with metric (4.16) describes the anti-self-dual form-field perturbations to the black hole in Taub-NUT space from the near horizon region perspective.
We also note that the three-dimensional part of the metric (4.16) is of the following special form,
@@
$ds^{2}=\ell^{2}\left(W^{2}dUdV+\frac{dW^{2}}{W^{2}}-W^{4}\bar{h}(\bar{v}(V))^{2}dV^{2}\right)+\ell^{2}d\Omega^{2}_{S^{3}}.$ (4.20)
@@ 5 Fermionic hair modes in the near horizon region
In this section, we make some observations on the non-normalisability of the fermionic hair modes of refs. [5, 4, 7] in $AdS_{3}\times S^{3}$. We find it easiest to present this analysis in six-dimensions. The relevant linearised equations for the gravitino $\psi^{\alpha}_{M}$ are
@@ 6 Discussion
    Using the AdS/CFT dictionary, we can identify the dimensions of the CFT operators that correspond to the bulk deformations discussed above. Turning on non-normalisable modes correspond to deformations of the CFT with operators of conformal dimension $\Delta>2$. For operators dual to scalars of mass $m$ in $AdS_{3}$, conformal dimension $\Delta$ is
    $\Delta=1+\sqrt{1+m^{2}\ell^{2}}.$ (6.1)
    The Garfinkle-Vachaspati deformations in $AdS_{3}\times S^{3}$ with $L=1$ and $L=2$ can be identified with scalars of masses $m^{2}\ell^{2}=L(L+2)$. Hence, the smooth hair mode with $L=2$ corresponds to a deformation of the CFT by an operator of mass dimensions
    $\Delta=1+\sqrt{1+8}=4.$ (6.2)
    The same conclusion can be reached by considering the Garfinkle-Vachaspati deformation as a metric deformation. Since the deformation term grows as $z^{-4}$ near the boundary $z=0$, it corresponds to a deformation of the CFT by an operator of mass dimensions $\Delta=4$, see, e.g., comments in [27].
@@
    We can similarly identify the dimensions of the operators that correspond to the form-field deformation. Taking one more exterior derivative of equation (4.7), we can convert it into the equation for a massive vector field
@@
    In section 4, we made the observation that the three-dimensional part of the metric (4.16) can be written in a special form (4.19). For the case when the function $h(v)$ is a constant, we recognise metric (4.19) as the null warped $AdS_{3}$ metric (and also as Schrödinger metric with $z=2$). As is well known, these metrics can be obtained as solutions to three-dimensional Einstein gravity with negative cosmological constant coupled to a massive vector field [28]. As discussed in the previous remark, the metric deformation discussed in section 4 can also be thought of as a deformation sourced by a massive vector field (6.3). Thus, the fact that the metric (4.16) can be written in the null warped $AdS_{3}$ form is perhaps not unexpected. Null warped $AdS_{3}$ metrics are known to be duals of deformations of the boundary CFT by operators of dimension $4$ [29]. This is in line with the observations made above. It will be interesting to identify such operators in the D1-D5 CFT and explore the connection to the null warped $AdS_{3}$ further.
In refs. [30, 31, 32] various deformations of $AdS_{3}\times S^{3}$ were considered in the context of the fuzzball program. Even though the starting point is the same, there are many differences between our results and their analyses. Firstly, since there are no horizons in the fuzzball cases, these references did not need to consider the regularity at the horizon. The analyses in these references is instead in global AdS. Secondly, the metric perturbations considered in these references do not have direct analogs in our analysis, but roughly speaking a class of modes considered in [30] corresponds to our $L=0$ Garfinkle-Vachaspati modes. [^7] See also point 3 on page 26 of ref. [30] where a related observation is made.
In ref. [33], it was suggested that fuzzballs should also admit hair modes like the ones that we have considered in this paper. Specifically, by putting fuzzballs in Taub-NUT space, it was conjectured that there should be fuzzball hair, like the ones obtained by putting the BMPV black hole in Taub-NUT space. We hope to report our progress on constructing such hair on a class of simplest fuzzballs in the future.